OpenStudy (cwrw238):

Differentiate sinsinsin2x

OpenStudy (cwrw238):

i'm confused with this one i guess you use the chain rule?

OpenStudy (anonymous):

You do, it's horrifying! :(

OpenStudy (anonymous):

d(sin(sin(sin(2x))))/dx=dsinu/du*du/dx u=sin(sin(2x)) and dsinu/du=cosu

OpenStudy (anonymous):

And then you apply it once more but that gets too messy for me to type - bound to make a mistake. Hehe, hope my answer helped to some extent atleast; good luck!

OpenStudy (anonymous):

*twice more

OpenStudy (anonymous):

Using the chain rule three times.

OpenStudy (cwrw238):

ugh - yes! ill try that

sam (.sam.):

you should be getting \[2 \cos (2 x) \cos (\sin (2 x)) \cos (\sin (\sin (2 x)))\]

OpenStudy (cwrw238):

i got 2cos(sinsin2x)*cos(sin2x)*2cos2x

OpenStudy (cwrw238):

not quite i've got one extra '2'

OpenStudy (anonymous):

Oh right, missed that one, sorry

OpenStudy (cwrw238):

i think sam is right though

OpenStudy (cwrw238):

yea the first 2 is wrong

OpenStudy (cwrw238):

thnx guys

OpenStudy (cwrw238):

i did it in steps from left to right

OpenStudy (anonymous):

Sam's correct! You should take derivative from RIGHT to left: ( sin2x )' = 2 cosx sin ( sin2x) ' = 2 cosx * cos ( sin2x ) sin [ sin ( sin2x)] ' = 2 cosx * cos ( sin2x ) * cos [ sin ( sin2x)] Just go step by step, you won't miss it :)